nolegirl24:

Write x^2 − 8x + 13 = 0 in the form (x − a)^2 = b, where a and b are integers. (x − 4)2 = 3 (x − 3)2 = 2 (x − 2)2 = 1 (x − 1)2 = 4

nolegirl24:

@jhonyy9

jhonyy9:

do you know formula (a-b)^2 = ?

nolegirl24:

yeah i think so

jhonyy9:

write it here please

nolegirl24:

yeah one sec

jhonyy9:

courage just like a first step

butterbear:

ya that will help

nolegirl24:

Would it be like this? (x-1)^a = 8

jhonyy9:

no

nolegirl24:

damnit ik where i messed up

jhonyy9:

you need to know formula (a-b)^2 = ?

nolegirl24:

the a would be 2

nolegirl24:

so nvm then ig idk it

butterbear:

ya i think so

jhonyy9:

a=2 not is right

butterbear:

hold on

nolegirl24:

it would be like this then cause i messed up (x-1)^2 = 8 right.....????????

jhonyy9:

and so how will get the 13 ?

jhonyy9:

ATTENTION please !

nolegirl24:

I am paying attention!!!!!! i was looking back at the damn problem!!!!

jhonyy9:

rewrite it in this way x^2 - 2*4x +13+3-3 = 0 bc. x^2 -8x +16 = (x-4)^2

jhonyy9:

from this result a=4 and so b = ?

nolegirl24:

I am just confused now this is retarded

jhonyy9:

ok. looke please (a-b)^2 = a^2 -2ab +b^2 yes ?

nolegirl24:

ig

jhonyy9:

so fro you can making from this x^2 -8x +13 perfect square of (x-4)^2 = x^2 -2*4x +16 what you need to do with 13 ?

jhonyy9:

you need adding to 13 a 3

nolegirl24:

Iḿma be honest i am completely confused and this is everywhere

jhonyy9:

but for you can adding this 3 you need subtracting it too for this expression not being modified

nolegirl24:

idk what to do and not to be rude but you are making this problem go everywhere

jhonyy9:

first of all you need making from x^2 -8x +13 a square perfect yes ?

jhonyy9:

this will be (x-a)^2

nolegirl24:

ok

jhonyy9:

you can rewriting this x^2 -8x +13 in this form x^2 -2*4x +13 yes ?

jhonyy9:

and bc. formula (a-b)^2 = a^2 -2ab +b^2 this mean that in place of b^2 you need getting 16 what is the square of 4

jhonyy9:

@Ultrilliam your opinion please

jhonyy9:

why dont wan believing me the asker that this is right ?

jhonyy9:

@Elsa213

jhonyy9:

@dude please

dude:

I don't think they understand you, I understand your goal and what you try to say but they don't I can try to help you out in terms of communicating with others, are you a native English speaker?

nolegirl24:

I got the answer

dude:

Okay

jhonyy9:

no my mather language is hungarian and second is roumanian bc. my birth country is Roumania but i live in Hungary

nolegirl24:

its (x-4)^2 = 3

jhonyy9:

yes exactly

jhonyy9:

thank you @dude